The cue ball moving horizontally at 6.2 m/s collides with stationary billiard ball. As a result, the billiard ball is deflected 45° from the horizontal. Find the speeds of the two balls and the angle the cue ball makes with the horizontal. Assume the collision is elastic and that the masses of the balls are equal.


1
Expert's answer
2021-03-23T11:08:38-0400



"v\cos45°=u'\to u'=6.5\cdot\cos45°=4.596\ (m/s)"


"v\sin45°=v'\to v'=6.5\cdot\sin45°=4.596\ (m/s)"


"\alpha=45°"